# Topotactical hydrogen induced single-band -wave superconductivity in La2NiO4

### v1

@@ Abstract
Figure 4(a) displays the DA Fermi surface at various hole dopings (10% , 15% and 20%) and temperatures (46 K and 116 K, based on the constructed 2D Hubbard model with the hopping parameters from Table 1. The Fermi surface is obtained from the lowest Matsubara frequency. Please note that the inter-layer (-direction) hopping between orbital is justified by the spacing layers, which makes H-La2NiO4 an even more ideal two-dimensional system than infinite-layer Nd(La)NiO2 (which has a -hopping of 34 meV compared to 2 meV in H-La2NiO4). For a doping level =0.85 (corresponding to H-La1.85Sr0.15NiO4), a well-defined hole-like Fermi surface can be observed in Figure 3(a). In contrast, at =0.9 (H-La1.9Sr0.1NiO4), strong AFM spin fluctuations induce significant scattering at the anti-nodal point =(,0). This signals the vicinity of the pseudogap regime. At the lower temperature of 46 K, the bands become more well-defined across the Brillouin zone, indicating a suppression of the hole contributions to, e.g., Hall coefficient at higher temperatures.
@@
The superconducting critical temperature can be determined from the divergence of the superconducting susceptibility , or equivalently, the leading superconducting eigenvalue . In matrix notation Kitatani et al. (2019), =/[1-], where is the bare superconducting susceptibility and is the irreducible vertex in the particle-particle channel calculated by DA. When the leading eigenvalue of approaches 1, the superconducting susceptibility diverges. Figure 4(b) plots versus temperature at different fillings of the Ni- orbital. It shows that approaches 1 at low temperatures for =0.80 and =0.85, whereas for =0.90 there seems to be no superconductivity. This indicates the absence of superconductivity at =0.90 at 13 K. This suggests a superconducting dome within the range 0.90n0.80. The momentum structure of the leading eigenvector (not shown) indicates -wave superconductivity.
@@
Conclusion. The recent discovery of infinite-layer nickelate superconductors has revealed a wide range of crystal structures, chemical compositions, and potential superconducting mechanisms in different nickelates. In contrast, although La2NiO4 has been extensively studied and exhibits a variety of novel quantum states, superconductivity has not been found yet experimentally. Here, we investigated the structural, thermal, magnetic, electronic, and superconducting properties of H-La2NiO4 where hydrogen is topotactically intercalated, using state-of-the-art theoretical methods. These methods include DA which is well-suited for calculating superconducting temperatures in two-dimensional strongly correlated oxides such as cuprates and nickelates. Unlike in LaNiO2H, it acts as an electronic donor in H-La2NiO4. This leads to a Ni+ valence with an electronic configuration similar to Cu2+. Also the in-plane AFM Mott insulating state for full hydrogen intercalation resembles that of cuprate superconductors such as La2CuO4. With an optimal hole doping of approximately 15%, our simulations predict that H-La2NiO4 exhibits a -wave superconducting state with up to 20 K, akin to its infinite-layer nickelate analogs. Against the background of recent experimental progress to add hydrogen via ionic liquid gating in perovskite and RP-oxides, the experimental synthesis of the suggested of H-La2NiO4 appears possible and promising.
I Acknowledgments
Acknowledgements.
@@
We thank Prof. Pu Yu, Viktor Christiansson and Eric Jacob for helpful insightful discussions. L. S. acknowledges support from the National Natural Science Foundation of China (Grant No. 12422407). L. S. and K. H. acknowledges funding through the Austrian Science Funds (FWF) project Grant DOI 10.55776/I5398. Calculations have been done on the the National Supercomputing Center (Xi’an) in Northwest University and the Vienna Scientific Cluster (VSC).References
- Norman (2016) M. Norman, Reports on Progress in Physics 79, 074502 (2016).
- Bednorz and Müller (1986) J. G. Bednorz and K. A. Müller, Zeitschrift für Physik B Condensed Matter 64, 189 (1986).
- Anderson (1997) P. W. Anderson, The theory of superconductivity in the high-Tc cuprate superconductors (Princeton University Press, 1997).
- Tsuei and Kirtley (2000) C. C. Tsuei and J. R. Kirtley, Rev. Mod. Phys. 72, 969 (2000).
- Damascelli et al. (2003) A. Damascelli, Z. Hussain, and Z.-X. Shen, Rev. Mod. Phys. 75, 473 (2003).
- Schilling et al. (1993) A. Schilling, M. Cantoni, J. Guo, and H. Ott, Nature 363, 56 (1993).
- Allan et al. (1989) N. Allan, J. Lawton, and W. Mackrodt, Philosophical Magazine B 59, 191 (1989).
- Bates and Eldridge (1989) F. E. Bates and J. Eldridge, Solid state communications 72, 187 (1989).
- Acrivos et al. (1994) J. Acrivos, M. C. Lei, C. Jiang, H. Nguyen, P. Metcalf, and J. Honig, Journal of Solid State Chemistry 111, 343 (1994).
- Lee and Pickett (2004) K.-W. Lee and W. E. Pickett, Phys. Rev. B 70, 165109 (2004).
- Anisimov et al. (1999) V. I. Anisimov, D. Bukhvalov, and T. M. Rice, Phys. Rev. B 59, 7901 (1999).
- Chaloupka and Khaliullin (2008) J. c. v. Chaloupka and G. Khaliullin, Phys. Rev. Lett. 100, 016404 (2008).
- Hansmann et al. (2009) P. Hansmann, X. Yang, A. Toschi, G. Khaliullin, O. K. Andersen, and K. Held, Phys. Rev. Lett. 103, 016401 (2009).
- Guo et al. (2018) H. Guo, Z. Li, L. Zhao, Z. Hu, C. Chang, C.-Y. Kuo, W. Schmidt, A. Piovano, T. Pi, O. Sobolev, et al., Nature communications 9, 43 (2018).
- Zhou et al. (2018) G. Zhou, F. Jiang, J. Zang, Z. Quan, and X. Xu, ACS applied materials & interfaces 10, 1463 (2018).
- Li et al. (2019a) D. Li, K. Lee, B. Y. Wang, M. Osada, S. Crossley, H. R. Lee, Y. Cui, Y. Hikita, and H. Y. Hwang, Nature 572, 624 (2019a).
- Li et al. (2020a) D. Li, B. Y. Wang, K. Lee, S. P. Harvey, M. Osada, B. H. Goodge, L. F. Kourkoutis, and H. Y. Hwang, Phys. Rev. Lett. 125, 027001 (2020a).
- Zeng et al. (2022) S. Zeng, C. Li, L. E. Chow, Y. Cao, Z. Zhang, C. S. Tang, X. Yin, Z. S. Lim, J. Hu, P. Yang, et al., Science advances 8, eabl9927 (2022).
- Osada et al. (2021) M. Osada, B. Y. Wang, B. H. Goodge, S. P. Harvey, K. Lee, D. Li, L. F. Kourkoutis, and H. Y. Hwang, Advanced Materials 33, 2104083 (2021).
- Osada et al. (2020) M. Osada, B. Y. Wang, K. Lee, D. Li, and H. Y. Hwang, Physical Review Materials 4, 121801 (2020).
- Pan et al. (2022) G. A. Pan, D. Ferenc Segedin, H. LaBollita, Q. Song, E. M. Nica, B. H. Goodge, A. T. Pierce, S. Doyle, S. Novakov, D. Córdova Carrizales, et al., Nature materials 21, 160 (2022).
- Yan et al. (2025) X. Yan, H. Zheng, Y. Li, H. Cao, D. P. Phelan, H. Zheng, Z. Zhang, H. Hong, G. Wang, Y. Liu, et al., Science Advances 11, eado4572 (2025).
- Sun et al. (2023) H. Sun, M. Huo, X. Hu, J. Li, Z. Liu, Y. Han, L. Tang, Z. Mao, P. Yang, B. Wang, et al., Nature 621, 493 (2023).
- Hou et al. (2023) J. Hou, P.-T. Yang, Z.-Y. Liu, J.-Y. Li, P.-F. Shan, L. Ma, G. Wang, N.-N. Wang, H.-Z. Guo, J.-P. Sun, et al., Chinese Physics Letters 40, 117302 (2023).
- Wang et al. (2024a) N. Wang, G. Wang, X. Shen, J. Hou, J. Luo, X. Ma, H. Yang, L. Shi, J. Dou, J. Feng, et al., Nature 634, 579 (2024a).
- Zhu et al. (2024) Y. Zhu, D. Peng, E. Zhang, B. Pan, X. Chen, L. Chen, H. Ren, F. Liu, Y. Hao, N. Li, et al., Nature 631, 531 (2024).
- Li et al. (2024) Q. Li, Y.-J. Zhang, Z.-N. Xiang, Y. Zhang, X. Zhu, and H.-H. Wen, Chinese Physics Letters 41, 017401 (2024).
- Pei et al. (2024) C. Pei, M. Zhang, D. Peng, S. Huangfu, S. Zhu, Q. Wang, J. Wu, Z. Xing, L. Zhang, Y. Chen, et al., arXiv preprint arXiv:2411.08677 (2024).
- Zhang et al. (2025) E. Zhang, D. Peng, Y. Zhu, L. Chen, B. Cui, X. Wang, W. Wang, Q. Zeng, and J. Zhao, arXiv preprint arXiv:2501.17709 (2025).
- Huang et al. (2024) X. Huang, H. Zhang, J. Li, M. Huo, J. Chen, Z. Qiu, P. Ma, C. Huang, H. Sun, and M. Wang, Chinese Physics Letters 41, 127403 (2024).
- Ko et al. (2024) E. K. Ko, Y. Yu, Y. Liu, L. Bhatt, J. Li, V. Thampy, C.-T. Kuo, B. Y. Wang, Y. Lee, K. Lee, et al., Nature , 1 (2024).
- Zhou et al. (2024) G. Zhou, W. Lv, H. Wang, Z. Nie, Y. Chen, Y. Li, H. Huang, W. Chen, Y. Sun, Q.-K. Xue, et al., arXiv preprint arXiv:2412.16622 (2024).
- Li et al. (2025a) P. Li, G. Zhou, W. Lv, Y. Li, C. Yue, H. Huang, L. Xu, J. Shen, Y. Miao, W. Song, et al., arXiv preprint arXiv:2501.09255 (2025a).
- Kitatani et al. (2023) M. Kitatani, L. Si, P. Worm, J. M. Tomczak, R. Arita, and K. Held, Phys. Rev. Lett. 130, 166002 (2023).
- Kitatani et al. (2020) M. Kitatani, L. Si, O. Janson, R. Arita, Z. Zhong, and K. Held, npj Quantum Materials 5, 59 (2020).
- Lechermann (2020a) F. Lechermann, Phys. Rev. B 101, 081110 (2020a).
- Lechermann (2020b) F. Lechermann, Phys. Rev. X 10, 041002 (2020b).
- Zhang et al. (2020a) G.-M. Zhang, Y.-f. Yang, and F.-C. Zhang, Phys. Rev. B 101, 020501 (2020a).
- Werner and Hoshino (2020) P. Werner and S. Hoshino, Phys. Rev. B 101, 041104 (2020).
- Zhang et al. (2020b) H. Zhang, L. Jin, S. Wang, B. Xi, X. Shi, F. Ye, and J.-W. Mei, Phys. Rev. Res. 2, 013214 (2020b).
- Sakakibara et al. (2020) H. Sakakibara, H. Usui, K. Suzuki, T. Kotani, H. Aoki, and K. Kuroki, Phys. Rev. Lett. 125, 077003 (2020).
- Zhang and Vishwanath (2020) Y.-H. Zhang and A. Vishwanath, Phys. Rev. Res. 2, 023112 (2020).
- Nomura et al. (2019) Y. Nomura, M. Hirayama, T. Tadano, Y. Yoshimoto, K. Nakamura, and R. Arita, Phys. Rev. B 100, 205138 (2019).
- Zhou et al. (2020) T. Zhou, Y. Gao, and Z. Wang, Science China Physics, Mechanics & Astronomy 63, 287412 (2020).
- Xie et al. (2022) T. Y. Xie, Z. Liu, C. Cao, Z. F. Wang, J. L. Yang, and W. Zhu, Phys. Rev. B 106, 035111 (2022).
- Li and Louie (2024) Z. Li and S. G. Louie, Phys. Rev. Lett. 133, 126401 (2024).
- Worm et al. (2022a) P. Worm, L. Si, M. Kitatani, R. Arita, J. M. Tomczak, and K. Held, Phys. Rev. Mater. 6, L091801 (2022a).
- Lechermann (2022) F. Lechermann, Phys. Rev. B 105, 155109 (2022).
- LaBollita and Botana (2022) H. LaBollita and A. S. Botana, Phys. Rev. B 105, 085118 (2022).
- LaBollita et al. (2022) H. LaBollita, M.-C. Jung, and A. S. Botana, Phys. Rev. B 106, 115132 (2022).
- Carrasco Álvarez et al. (2024) A. A. Carrasco Álvarez, S. Petit, L. Iglesias, M. Bibes, W. Prellier, and J. Varignon, Phys. Rev. B 110, 235107 (2024).
- Craco et al. (2022) L. Craco, A. S. de Arruda, and S. Leoni, Phys. Rev. Res. 4, 043036 (2022).
- Cheng et al. (2024) B. Cheng, D. Cheng, K. Lee, L. Luo, Z. Chen, Y. Lee, B. Y. Wang, M. Mootz, I. E. Perakis, Z.-X. Shen, et al., Nature Materials 23, 775 (2024).
- Cheng et al. (2025) B. Cheng, D. Cheng, K. Lee, M. Mootz, C. Huang, L. Luo, Z. Chen, Y. Lee, B. Y. Wang, I. E. Perakis, Z.-X. Shen, H. Y. Hwang, and J. Wang, Phys. Rev. B 111, 014519 (2025).
- Worm et al. (2024) P. Worm, Q. Wang, M. Kitatani, I. Biało, Q. Gao, X. Ren, J. Choi, D. Csontosová, K.-J. Zhou, X. Zhou, Z. Zhu, L. Si, J. Chang, J. M. Tomczak, and K. Held, Phys. Rev. B 109, 235126 (2024).
- Zhang et al. (2024a) Y. Zhang, L.-F. Lin, A. Moreo, T. A. Maier, and E. Dagotto, Nature Communications 15, 2470 (2024a).
- Yang et al. (2024) Q.-G. Yang, K.-Y. Jiang, D. Wang, H.-Y. Lu, and Q.-H. Wang, Phys. Rev. B 109, L220506 (2024).
- Zhang et al. (2023) Y. Zhang, L.-F. Lin, A. Moreo, T. A. Maier, and E. Dagotto, Phys. Rev. B 108, 165141 (2023).
- Zhang et al. (2024b) Y. Zhang, L.-F. Lin, A. Moreo, T. A. Maier, and E. Dagotto, Phys. Rev. Lett. 133, 136001 (2024b).
- Yang et al. (2023) Q.-G. Yang, D. Wang, and Q.-H. Wang, Phys. Rev. B 108, L140505 (2023).
- Zhang et al. (2024c) M. Zhang, H. Sun, Y.-B. Liu, Q. Liu, W.-Q. Chen, and F. Yang, Phys. Rev. B 110, L180501 (2024c).
- Liu et al. (2023) Y.-B. Liu, J.-W. Mei, F. Ye, W.-Q. Chen, and F. Yang, Phys. Rev. Lett. 131, 236002 (2023).
- Ni et al. (2025) X.-S. Ni, Y. Ji, L. He, T. Xie, D.-X. Yao, M. Wang, and K. Cao, npj Quantum Materials 10, 17 (2025).
- LaBollita et al. (2024) H. LaBollita, V. Pardo, M. R. Norman, and A. S. Botana, Phys. Rev. Mater. 8, L111801 (2024).
- Li et al. (2025b) Y. Li, Y. Cao, L. Liu, P. Peng, H. Lin, C. Pei, M. Zhang, H. Wu, X. Du, W. Zhao, et al., Science Bulletin 70, 180 (2025b).
- Chen et al. (2024) K. Chen, X. Liu, J. Jiao, M. Zou, C. Jiang, X. Li, Y. Luo, Q. Wu, N. Zhang, Y. Guo, and L. Shu, Phys. Rev. Lett. 132, 256503 (2024).
- Samarakoon et al. (2023) A. M. Samarakoon, J. Strempfer, J. Zhang, F. Ye, Y. Qiu, J.-W. Kim, H. Zheng, S. Rosenkranz, M. R. Norman, J. F. Mitchell, and D. Phelan, Phys. Rev. X 13, 041018 (2023).
- Yi et al. (2024) X.-W. Yi, Y. Meng, J.-W. Li, Z.-W. Liao, W. Li, J.-Y. You, B. Gu, and G. Su, Phys. Rev. B 110, L140508 (2024).
- Meng et al. (2024) Y. Meng, Y. Yang, H. Sun, S. Zhang, J. Luo, L. Chen, X. Ma, M. Wang, F. Hong, X. Wang, et al., Nature Communications 15, 10408 (2024).
- Kumar et al. (2020) S. Kumar, Ø. Fjellvåg, A. O. Sjåstad, and H. Fjellvåg, Journal of Magnetism and Magnetic Materials 496, 165915 (2020).
- Liu et al. (2024) Z. Liu, M. Huo, J. Li, Q. Li, Y. Liu, Y. Dai, X. Zhou, J. Hao, Y. Lu, M. Wang, et al., Nature Communications 15, 7570 (2024).
- Leonov (2024) I. V. Leonov, Phys. Rev. B 109, 235123 (2024).
- Lander et al. (1989) G. H. Lander, P. J. Brown, J. Spal/ek, and J. M. Honig, Phys. Rev. B 40, 4463 (1989).
- Pintschovius et al. (1989) L. Pintschovius, J. M. Bassat, P. Odier, F. Gervais, G. Chevrier, W. Reichardt, and F. Gompf, Phys. Rev. B 40, 2229 (1989).
- Brown (1992) I. Brown, Zeitschrift für Kristallographie 199, 255 (1992).
- Rodriguez-Carvajal et al. (1991) J. Rodriguez-Carvajal, M. Fernandez-Diaz, and J. Martinez, Journal of Physics: Condensed Matter 3, 3215 (1991).
- Wang et al. (2024b) X. Wang, P. Worm, Y. Gao, W. Wu, N. Liu, Y. Wang, K. Held, and L. Si, Phys. Rev. B 110, 205110 (2024b).
- Biało et al. (2024) I. Biało, L. Martinelli, G. De Luca, P. Worm, A. Drewanowski, S. Jöhr, J. Choi, M. Garcia-Fernandez, S. Agrestini, K.-J. Zhou, et al., Communications Physics 7, 230 (2024).
- Cui et al. (2024) T. Cui, S. Choi, T. Lin, C. Liu, G. Wang, N. Wang, S. Chen, H. Hong, D. Rong, Q. Wang, et al., Communications Materials 5, 32 (2024).
- Guo and Temmerman (1988) G. Guo and W. Temmerman, Journal of Physics C: Solid State Physics 21, L803 (1988).
- Mahadevan et al. (1997) P. Mahadevan, K. Sheshadri, D. D. Sarma, H. R. Krishnamurthy, and R. Pandit, Phys. Rev. B 55, 9203 (1997).
- Homes et al. (2007) C. C. Homes, J. M. Tranquada, and D. J. Buttrey, Phys. Rev. B 75, 045128 (2007).
- Hücker et al. (2004) M. Hücker, K. Chung, M. Chand, T. Vogt, J. M. Tranquada, and D. J. Buttrey, Phys. Rev. B 70, 064105 (2004).
- Jestädt et al. (1999) T. Jestädt, K. H. Chow, S. J. Blundell, W. Hayes, F. L. Pratt, B. W. Lovett, M. A. Green, J. E. Millburn, and M. J. Rosseinsky, Phys. Rev. B 59, 3775 (1999).
- Jung et al. (2001) J. H. Jung, D.-W. Kim, T. W. Noh, H. C. Kim, H.-C. Ri, S. J. Levett, M. R. Lees, D. M. Paul, and G. Balakrishnan, Phys. Rev. B 64, 165106 (2001).
- Scalapino (1995) D. J. Scalapino, Physics Reports 250, 329 (1995).
- Li et al. (2020b) Z. Li, S. Shen, Z. Tian, K. Hwangbo, M. Wang, Y. Wang, F. M. Bartram, L. He, Y. Lyu, Y. Dong, et al., Nature Communications 11, 184 (2020b).
- Li et al. (2019b) H.-B. Li, F. Lou, Y. Wang, Y. Zhang, Q. Zhang, D. Wu, Z. Li, M. Wang, T. Huang, Y. Lyu, et al., Advanced Science 6, 1901432 (2019b).
- Lu et al. (2022) N. Lu, Z. Zhang, Y. Wang, H.-B. Li, S. Qiao, B. Zhao, Q. He, S. Lu, C. Li, Y. Wu, et al., Nature Energy 7, 1208 (2022).
- Shen et al. (2021) S. Shen, Z. Li, Z. Tian, W. Luo, S. Okamoto, and P. Yu, Phys. Rev. X 11, 021018 (2021).
- Tian et al. (2024) D. Tian, L. Miao, L. Si, N. J. Schreiber, S. Shen, J. Zhang, X. Shu, X. Wang, H. P. Nair, J. P. Ruf, D. G. Schlom, K. M. Shen, and P. Yu, Phys. Rev. Mater. 8, 074408 (2024).
- Hohenberg and Kohn (1964) P. Hohenberg and W. Kohn, Phys. Rev. 136, B864 (1964).
- Kohn and Sham (1965) W. Kohn and L. J. Sham, Phys. Rev. 140, A1133 (1965).
- Metzner and Vollhardt (1989) W. Metzner and D. Vollhardt, Phys. Rev. Lett. 62, 324 (1989).
- Georges et al. (1996) A. Georges, G. Kotliar, W. Krauth, and M. J. Rozenberg, Rev. Mod. Phys. 68, 13 (1996).
- Kotliar and Vollhardt (2004) G. Kotliar and D. Vollhardt, Physics Today 57, 53 (2004).
- Held (2007) K. Held, Advances in physics 56, 829 (2007).
- Held et al. (2008) K. Held, A. A. Katanin, and A. Toschi, Progress of Theoretical Physics Supplement 176, 117 (2008).
- Galler et al. (2019) A. Galler, P. Thunström, J. Kaufmann, M. Pickem, J. M. Tomczak, and K. Held, Computer Physics Communications 245, 106847 (2019).
- Rohringer et al. (2018) G. Rohringer, H. Hafermann, A. Toschi, A. A. Katanin, A. E. Antipov, M. I. Katsnelson, A. I. Lichtenstein, A. N. Rubtsov, and K. Held, Rev. Mod. Phys. 90, 025003 (2018).
- Kresse and Furthmüller (1996) G. Kresse and J. Furthmüller, Computational materials science 6, 15 (1996).
- Kresse and Furthmüller (1996) G. Kresse and J. Furthmüller, Phys. Rev. B 54, 11169 (1996).
- Blaha et al. (2001) P. Blaha, K. Schwarz, G. Madsen, D. Kvasnicka, and J. Luitz, Wien2k,an augmented plane wave+ local orbitals program for calculating crystal properties (2001).
- Schwarz et al. (2002) K. Schwarz, P. Blaha, and G. K. H. Madsen, Comp. Phys. Comm. 147, 71 (2002).
- Perdew et al. (1996) J. P. Perdew, K. Burke, and M. Ernzerhof, Phys. Rev. Lett. 77, 3865 (1996).
- Miyake and Aryasetiawan (2008) T. Miyake and F. Aryasetiawan, Phys. Rev. B 77, 085122 (2008).
- Si et al. (2025) L. Si, P. Liu, and C. Franchini, Phys. Rev. Mater. 9, 015001 (2025).
- Kresse and Hafner (1993a) G. Kresse and J. Hafner, Phys. Rev. B 48, 13115 (1993a).
- Kresse and Hafner (1994) G. Kresse and J. Hafner, Phys. Rev. B 49, 14251 (1994).
- Kresse and Hafner (1993b) G. Kresse and J. Hafner, Phys. Rev. B 47, 558 (1993b).
- Wannier (1937) G. H. Wannier, Phys. Rev. 52, 191 (1937).
- Marzari et al. (2012) N. Marzari, A. A. Mostofi, J. R. Yates, I. Souza, and D. Vanderbilt, Rev. Mod. Phys. 84, 1419 (2012).
- Mostofi et al. (2008) A. A. Mostofi, J. R. Yates, Y.-S. Lee, I. Souza, D. Vanderbilt, and N. Marzari, Computer Physics Communications 178, 685 (2008).
- Kuneš et al. (2010) J. Kuneš, R. Arita, P. Wissgott, A. Toschi, H. Ikeda, and K. Held, Computer Physics Communications 181, 1888 (2010).
- Liechtenstein et al. (1995) A. I. Liechtenstein, V. I. Anisimov, and J. Zaanen, Phys. Rev. B 52, R5467 (1995).
- Anisimov et al. (1993) V. I. Anisimov, I. V. Solovyev, M. A. Korotin, M. T. Czyżyk, and G. A. Sawatzky, Phys. Rev. B 48, 16929 (1993).
- Gull et al. (2011) E. Gull, A. J. Millis, A. I. Lichtenstein, A. N. Rubtsov, M. Troyer, and P. Werner, Rev. Mod. Phys. 83, 349 (2011).
- Parragh et al. (2012) N. Parragh, A. Toschi, K. Held, and G. Sangiovanni, Phys. Rev. B 86, 155158 (2012).
- Wallerberger et al. (2019) M. Wallerberger, A. Hausoel, P. Gunacker, A. Kowalski, N. Parragh, F. Goth, K. Held, and G. Sangiovanni, Computer Physics Communications 235, 388 (2019).
- Gubernatis et al. (1991) J. E. Gubernatis, M. Jarrell, R. N. Silver, and D. S. Sivia, Phys. Rev. B 44, 6011 (1991).
- Sandvik (1998) A. W. Sandvik, Phys. Rev. B 57, 10287 (1998).
- Kaufmann and Held (2023) J. Kaufmann and K. Held, Computer Physics Communications 282, 108519 (2023).
- Kitatani et al. (2019) M. Kitatani, T. Schäfer, H. Aoki, and K. Held, Phys. Rev. B 99, 041115 (2019).
- Kitatani et al. (2022) M. Kitatani, R. Arita, T. Schäfer, and K. Held, Journal of Physics: Materials 5, 034005 (2022).
- Si et al. (2020) L. Si, W. Xiao, J. Kaufmann, J. M. Tomczak, Y. Lu, Z. Zhong, and K. Held, Phys. Rev. Lett. 124, 166402 (2020).
- Lee et al. (2023) K. Lee, B. Y. Wang, M. Osada, B. H. Goodge, T. C. Wang, Y. Lee, S. Harvey, W. J. Kim, Y. Yu, C. Murthy, et al., Nature 619, 288 (2023).
- Note (1) In this study, we initially designed several possible structures of H-La2NiO4 by intercalating hydrogen into La2NiO4, guided by crystal symmetry analysis. Subsequent non-magnetic DFT calculations were performed to optimize these structures and determine the ground state of H-La2NiO4. However, a more comprehensive structural search may require exploring a broader range of possible configurations within larger crystal cells. Our primary focus in this research is to investigate the electron doping effect introduced by hydrogen intercalation in La2NiO4.
- (128) Supplemental information regarding additional results and computational details of DFT, DFT+ and AIMD calculations, as well as spectra, are available at XXX.
- Tran and Blaha (2009) F. Tran and P. Blaha, Phys. Rev. Lett. 102, 226401 (2009).
- Si et al. (2023) L. Si, P. Worm, D. Chen, and K. Held, Phys. Rev. B 107, 165116 (2023).
- Han et al. (2023) X. Han, P. Li, X. Sui, H. Jin, F. Luo, L. Qiao, and B. Huang, Phys. Rev. B 108, 165145 (2023).
- Marzari and Vanderbilt (1997) N. Marzari and D. Vanderbilt, Phys. Rev. B 56, 12847 (1997).
- Souza et al. (2001) I. Souza, N. Marzari, and D. Vanderbilt, Phys. Rev. B 65, 035109 (2001).
- Uemura et al. (1989) Y. J. Uemura, G. M. Luke, B. J. Sternlieb, J. H. Brewer, J. F. Carolan, W. N. Hardy, R. Kadono, J. R. Kempton, R. F. Kiefl, S. R. Kreitzman, P. Mulhern, T. M. Riseman, D. L. Williams, B. X. Yang, S. Uchida, H. Takagi, J. Gopalakrishnan, A. W. Sleight, M. A. Subramanian, C. L. Chien, M. Z. Cieplak, G. Xiao, V. Y. Lee, B. W. Statt, C. E. Stronach, W. J. Kossler, and X. H. Yu, Phys. Rev. Lett. 62, 2317 (1989).
- Padilla et al. (2005) W. J. Padilla, Y. S. Lee, M. Dumm, G. Blumberg, S. Ono, K. Segawa, S. Komiya, Y. Ando, and D. N. Basov, Phys. Rev. B 72, 060511 (2005).
- Worm et al. (2022b) P. Worm, M. Kitatani, J. M. Tomczak, L. Si, and K. Held, Phys. Rev. B 105, 085110 (2022b).

### v2

@@ Abstract
Figure 4(a) displays the DA Fermi surface at various hole dopings (10%, 15% and 20%) and temperatures (46 K and 116 K), based on the constructed 2D Hubbard model with the hopping parameters from Table 1. The Fermi surface is obtained from the lowest Matsubara frequency. Please note that the interlayer (-direction) hopping between orbital is justified by the spacing layers, which makes H-La2NiO4 an even more ideal two-dimensional system than infinite-layer Nd(La)NiO2 (which has a hopping of 34 meV compared to 2 meV in H-La2NiO4). For a doping level =0.85 (corresponding to H-La1.85Sr0.15NiO4), a well-defined holelike Fermi surface can be observed in Fig. 4(a). In contrast, at =0.9 (H-La1.9Sr0.1NiO4), strong AFM spin fluctuations induce significant scattering at the antinodal point =(,0). This signals the vicinity of the pseudogap regime. At the lower temperature of 46 K, the bands become more well-defined across the Brillouin zone, indicating a suppression of the hole contributions to, e.g., Hall coefficient at higher temperatures.
@@
The superconducting critical temperature can be determined from the divergence of the superconducting susceptibility , or equivalently, the leading superconducting eigenvalue . In matrix notation Kitatani et al. (2019), =/[1-], where is the bare superconducting susceptibility and is the irreducible vertex in the particle-particle channel calculated by DA. When the leading eigenvalue of approaches 1, the superconducting susceptibility diverges. Figure 4(b) plots versus temperature at different fillings of the Ni- orbital. It shows that approaches 1 at low temperatures for =0.80 and =0.85, whereas for =0.90, there seems to be no superconductivity. Specifically, the superconducting state is absent for =0.9 down to the lowest measured temperature, and for =0.8, it is only observed at T13 K. This suggests a superconducting dome within the range 0.90n0.80. The momentum structure of the leading eigenvector (not shown) indicates -wave superconductivity.
@@
Conclusion—The recent discovery of infinite-layer nickelate superconductors has revealed a wide range of crystal structures, chemical compositions, and potential superconducting mechanisms in different nickelates. In contrast, although La2NiO4 has been extensively studied and exhibits a variety of novel quantum states, superconductivity has not been found yet experimentally. Here, we investigated the structural, thermal, magnetic, electronic, and superconducting properties of H-La2NiO4 where hydrogen is topotactically intercalated, using state-of-the-art theoretical methods. These methods include DA, which is well-suited for calculating superconducting temperatures in two-dimensional strongly correlated oxides such as cuprates and nickelates. Unlike in LaNiO2H, it acts as an electronic donor in H-La2NiO4. This leads to a Ni+ valence with an electronic configuration similar to Cu2+. Also the in-plane AFM Mott insulating state for full hydrogen intercalation resembles that of cuprate superconductors such as La2CuO4. With an optimal hole doping of approximately 15%, our simulations predict that H-La2NiO4 exhibits a -wave superconducting state with up to 20 K, akin to its infinite-layer nickelate analogs. Against the background of recent experimental progress to add hydrogen via ionic liquid gating in perovskite and RP-oxides, the experimental synthesis of the suggested H-La2NiO4 appears possible and promising.
@@
Acknowledgments—We thank Prof. Pu Yu, Viktor Christiansson and Eric Jacob for helpful insightful discussions. L. S. acknowledges support from the National Natural Science Foundation of China (Grant No. 12422407). L. S. and K. H. acknowledge funding through the Austrian Science Funds (FWF) Project Grant DOI 10.55776/I5398. W.-F. W. acknowledges the funding from the China Scholarship Council (CSC). Calculations have been done on the the National Supercomputing Center (Xi’an) in Northwest University and the Vienna Scientific Cluster (VSC).